# Statistical Inference for Explainable Boosting Machines

### v1

@@ E.3.2 Same-feature Bias for Algorithm 1
By Lemma D.1 (a direct consequence of Assumption 4.7), on $\mathsf{H}_{0}$ we have
@@
E.3.3 Same-feature Bias for Algorithm 2
Lemma E.10 (Same–feature Bias for Algorithm 2).
@@ E.3.3 Same-feature Bias for Algorithm 2
E.3.4 Rate of Leakage for Algorithm 1
@@
Lemma E.11 (Rate of Leakage for Algorithm 1).
@@
By Lemma E.3, we know that $\mathbf{I}+\mathbf{J}_{n}\mathbf{K}$ is positive semidefinite, has bounded inverse and has bounded spectrum from $1+\lambda_{\text{min}}(\mathcal{K}),1+\lambda_{\text{max}}(\mathcal{K})]$. Hence by proposition F.2, the bound above proceeds as
@@
E.3.5 Rate of Leakage for Algorithm 2
@@
Lemma E.12 (Rate of Leakage for Algorithm 2).
@@
can be bound in the same manner, since E.2 guarantees $\left\|\mathbf{I}-\mathbf{K}^{(k)}\right\|_{\text{op}}\leq\frac{1}{1-\rho_{k}}$ and $\left\|\mathbf{I}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}\right\|_{\text{op}}\leq 1$. Hence the goal is to control the rate of the last linear transformation.
@@
By Lemma E.5, we know that $\mathbf{I}+\mathcal{K}$ is positive semidefinite, has bounded inverse and has bounded spectrum from $1+\lambda_{\text{min}}(\mathcal{K}),1+\lambda_{\text{max}}(\mathcal{K})]$. Hence by proposition F.2, the bound above proceeds as
@@
Collecting Lemmas E.8E.12 and the we conclude that the random-design bias terms vanish in probability. Together with the conditional CLT for the noise term, Slutsky’s theorem implies the full estimator is asymptotically normal (formalized in the next subsection).
E.4 Unconditional Asymptotic Normality
In Section E.3, we have shown that all bias terms vanish to 0 deterministically, given the rate on $\left\|\mathbf{k}^{(k)}\right\|$ and $\left\|\mathbf{r}^{(k)}\right\|$ hold. However the rate does not come for free. To extend them to a random design case, one would need a Kolmogorov extension theorem to define a new probability space and the infinte series $(\mathbf{x}_{i},\mathbf{y}_{i}),i=1,\cdots,\infty$’s projection onto finite series. See (Zhou and Hooker, 2019) for detailed construction. It can be shown that conditional asymptotic normality can be extended to the random design case. We will drop the subscript $A,B$ now since they share the same $O_{p}(\cdot)$ rate established in Section E.3.
Lemma E.13.
@@ E.4 Unconditional Asymptotic Normality
Combining all results from section E.3, E.2, E.4.
Theorem E.14 (Asymptotic Normality).
@@
Theorem E.15 (CLT for Baseline Vector).
@@
Appendix F Stochastic Contraction Mapping Theorem
Theorem F.1.
@@ Appendix F Stochastic Contraction Mapping Theorem
Proposition F.2 (Demko–Moss–Smith, General Sparse Exponential Decay).
@@
Appendix G Additional Figures
G.1 Study of Obesity Data (Palechor and la Hoz Manotas, 2019)
Figure 7: Full set of estimated feature effects for Algorithm 1 and EBM proper (Nori et al., 2019), when regressing weight on the given set of covariates. Algorithm 1 by-and-large yields qualitatively similar estimates, with confidence intervals for each term displayed in the shaded area. The terms roughly agree apart from the effect of gender on weight, which we argue should be higher for men than for women.

### v2

@@ F.3.2 Same-feature Bias for Algorithm 1
By Lemma E.1 (a direct consequence of Assumption 4.7), on $\mathsf{H}_{0}$ we have
@@
F.3.3 Same-feature Bias for Algorithm 2
Lemma F.10 (Same–feature Bias for Algorithm 2).
@@ F.3.3 Same-feature Bias for Algorithm 2
F.3.4 Rate of Leakage for Algorithm 1
@@
Lemma F.11 (Rate of Leakage for Algorithm 1).
@@
By Lemma F.3, we know that $\mathbf{I}+\mathbf{J}_{n}\mathbf{K}$ is positive semidefinite, has bounded inverse and has bounded spectrum from $1+\lambda_{\text{min}}(\mathcal{K}),1+\lambda_{\text{max}}(\mathcal{K})]$. Hence by proposition G.2, the bound above proceeds as
@@
F.3.5 Rate of Leakage for Algorithm 2
@@
Lemma F.12 (Rate of Leakage for Algorithm 2).
@@
can be bound in the same manner, since F.2 guarantees $\left\|\mathbf{I}-\mathbf{K}^{(k)}\right\|_{\text{op}}\leq\frac{1}{1-\rho_{k}}$ and $\left\|\mathbf{I}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}\right\|_{\text{op}}\leq 1$. Hence the goal is to control the rate of the last linear transformation.
@@
By Lemma F.5, we know that $\mathbf{I}+\mathcal{K}$ is positive semidefinite, has bounded inverse and has bounded spectrum from $1+\lambda_{\text{min}}(\mathcal{K}),1+\lambda_{\text{max}}(\mathcal{K})]$. Hence by proposition G.2, the bound above proceeds as
@@
Collecting Lemmas F.8F.12 and the we conclude that the random-design bias terms vanish in probability. Together with the conditional CLT for the noise term, Slutsky’s theorem implies the full estimator is asymptotically normal (formalized in the next subsection).
F.4 Unconditional Asymptotic Normality
In Section F.3, we have shown that all bias terms vanish to 0 deterministically, given the rate on $\left\|\mathbf{k}^{(k)}\right\|$ and $\left\|\mathbf{r}^{(k)}\right\|$ hold. However the rate does not come for free. To extend them to a random design case, one would need a Kolmogorov extension theorem to define a new probability space and the infinte series $(\mathbf{x}_{i},\mathbf{y}_{i}),i=1,\cdots,\infty$’s projection onto finite series. See Zhou and Hooker (2019) for detailed construction. It can be shown that conditional asymptotic normality can be extended to the random design case. We will drop the subscript $A,B$ now since they share the same $O_{p}(\cdot)$ rate established in Section F.3.
Lemma F.13.
@@ F.4 Unconditional Asymptotic Normality
Combining all results from section F.3, F.2, F.4.
Theorem F.14 (Asymptotic Normality).
@@
Theorem F.15 (CLT for Baseline Vector).
@@
Appendix G Stochastic Contraction Mapping Theorem
Theorem G.1.
@@ Appendix G Stochastic Contraction Mapping Theorem
Proposition G.2 (Demko–Moss–Smith, General Sparse Exponential Decay).
@@
Appendix H Additional Figures
H.1 Study of Obesity Data Palechor and la Hoz Manotas (2019)
Figure 7: Full set of estimated feature effects for Algorithm 1 and EBM proper Nori et al. (2019), when regressing weight on the given set of covariates. Algorithm 1 by-and-large yields qualitatively similar estimates, with confidence intervals for each term displayed in the shaded area. The terms roughly agree apart from the effect of gender on weight, which we argue should be higher for men than for women.